OpenStudy (anonymous):

Finding the exact value of Sin[cos^-1(0)]

OpenStudy (anonymous):

I know cos=0 At pi/2 and 3pi/2 but not sure where to go from there

OpenStudy (dtan5457):

cos=x sin=y cos=0 at 90 degrees and 270 degrees

OpenStudy (dtan5457):

so im pretty sure cos^-1 would be...

OpenStudy (dtan5457):

I forgot.

OpenStudy (michele_laino):

hint: \[\large \arccos 0 = \frac{\pi }{2},\;\frac{{3\pi }}{2}\]

OpenStudy (anonymous):

Already said I got that far :p

OpenStudy (dtan5457):

Are you suggesting to simply do sin x pi/2, etc?

OpenStudy (michele_laino):

hint: what is sin(pi/2)=...? and: sin(3*pi/2)=...?

OpenStudy (dtan5457):

OHH

OpenStudy (dtan5457):

You would convert those radians to degrees first, right? Then use the unit circle?

OpenStudy (anonymous):

Sin(pi/2)=0.02741 it's that simple?

OpenStudy (dtan5457):

Well, see that's the confusing part sin(90 degrees) would be 1.

OpenStudy (dtan5457):

depends on how we convert it. @Michele_Laino

OpenStudy (michele_laino):

we have: sin(pi/2)= sin(90) = 1 and: sin(3*pi/2)=sin(270)=-1

OpenStudy (dtan5457):

oh, so we do convert it to degrees. all clear now.

OpenStudy (anonymous):

Thanks

OpenStudy (michele_laino):

thanks!